# Immiscible Diffusion: Accelerating Diffusion Training with Noise Assignment

### v1

@@ Abstract
In this paper, we point out suboptimal noise-data mapping leads to slow training of diffusion models. During diffusion training, current methods diffuse each image across the entire noise space, resulting in a mixture of all images at every point in the noise layer. We emphasize that this random mixture of noise-data mapping complicates the optimization of the denoising function in diffusion models. Drawing inspiration from the immiscible phenomenon in physics, we propose Immiscible Diffusion, a simple and effective method to improve the random mixture of noise-data mapping. In physics, miscibility can vary according to various intermolecular forces. Thus, immiscibility means that the mixing of the molecular sources is distinguishable. Inspired by this concept, we propose an assignment-then-diffusion training strategy. Specifically, prior to diffusing the image data into noise, we assign diffusion target noise for the image data by minimizing the total image-noise pair distance in a mini-batch. The assignment functions analogously to external forces to separate the diffuse-able areas of images, thus mitigating the inherent difficulties in diffusion training. Our approach is remarkably simple, requiring only one line of code to restrict the diffuse-able area for each image while preserving the Gaussian distribution of noise. This ensures that each image is projected only to nearby noise. To address the high complexity of the assignment algorithm, we employs a quantized-assignment strategy, which significantly reduces the computational overhead to a negligible level (e.g., 22.8ms for a large batch size of 1024 on an A6000). Experiments demonstrate that our method can achieve up to 3x faster training for consistency models and DDIM on the CIFAR dataset, and up to 1.3x faster on CelebA datasets for consistency models. Besides, we conduct thorough analysis about the Immiscible Diffusion, which sheds lights on how it improves diffusion training speed while improving the fidelity. The code is available at https://yhli123.github.io/immiscible-diffusion
Figure 1: Immiscible Diffusion uses a single line of code to efficiently re-assign a batch of noise to images. This process results in only a 2% reduction in distance post-assignment, leading to up to 3x increased training efficiency while maintaining a reasonable FID on top of the Consistency Model for CIFAR Dataset. Additionally, Immiscible Diffusion significantly enhances the image quality of Stable Diffusion for both unconditional and conditional generation tasks on ImageNet Dataset within the same number of training steps.
@@ 1 Introduction
The diffusion model has made impressive progress in image generation by framing the process as a phase of denoising random Gaussian noise into the final image. Despite the advancements, training a diffusion model is resource-intensive. For example, even on the primary image dataset CIFAR-10, the representative few-step diffusion model, Consistency Model [38], requires training for 10 days on 4 A6000 GPUs to reach a desired FID score of around 10. Similarly, with fewer model parameters, multiple-step diffusion model DDIM [35] still requires 24 hours on an A5000 GPU on the CIFAR-10 dataset. Although recent remarkable achievements in accelerating the inference of diffusion models [13, 25, 38, 20, 22, 23] have been accomplished, the inefficiency of diffusion training remains a significant bottleneck, hindering the iterative development of vision generative AI.
@@
Previous methods to improve diffusion training have focused on various strategies, such as balancing the impact of activation layers and neural weights [11], modifying the hyperparameters and design choices [37], and leveraging patchifying strategies [43] etc. Specifically, Karras et al. [11] modifies the activation magnitude, neural weight standardization, and group normalization, achieving significant acceleration in diffusion training. Besides, previous work [37] proposes a customized method for consistency model to improve the performance and diffusion training. Our method is orthogonal to these previous methods. We got inspired by the Immiscible Diffusion in physics. As illustrated in Fig. 2, miscible particles tightly jumble together during the diffusion process, making it difficult to separate them individually during denoising phase. However, when the particles are rendered immiscible, they can still achieve a similar overall distribution while remaining clearly distinguishable (see the shape of the two images at the right of Fig. 2). This insight informs our strategy for improving the disentanglement of diffused data.
@@
We draw an analogy from the phenomenon of Immiscible Diffusion and relate the distribution of image data to the behavior of particles discussed above. In traditional diffusion processes, each image can be diffused to any point in the noise space, and conversely, each point in the noise space can be denoised to any source image, as illustrated in the left image of Fig. 2 (b). We hypothesize that the jumbled image-noise mapping creates a miscible diffusion effect and makes the optimization of diffusion model difficult. Inspired by the Immiscible Diffusion, we are motivated to make the mixed diffusion phase distinguishable.
@@
Figure 2: Physical illustration of Immiscible Diffusion. (a) depict the miscible and Immiscible Diffusion phenomenon in physics, while (b) demonstrate the image-noise pair relation in traditional (miscible) and Immiscible Diffusion method.
@@
We propose a simple and novel Immiscible Diffusion method. Note that we still sample Gaussian noise but perform a batch-wise assignment of noise to each image based on the distance between them during training. This approach ensures that each image is only diffused to surrounding areas while maintaining the overall Gaussian distribution of all noises. This noise assignment makes the diffusion model especially effective at denoising at high noise levels, which benefits the current trend of few-step denoising models. However, technically, the image data-noise assignment is an O($N^{2}logN$)-O($N^{3}$) operation, which introduces significant overhead during training, especially for large-scale training with huge batch sizes and high-resolution images. To address this, we employ a novel quantization method during assignment. We quantize the noise and image data into low-precision formats (e.g., 16-bit) during conducting the assignment algorithm. We highlight our method only involves one line of code, and is only performed during the training phase without modifying the model architecture, the noise scheduler, the sampler, or the method of inference.
@@
We conduct experiments on three diffusion baselines, Consistency models, DDIM and Stable Diffusion on the CIFAR-10, CelebA and ImageNet datasets. Results show that our proposed method significantly improves the training efficiency in all experiments. Specifically, we achieve 3x training efficiency for CIFAR-10 dataset with immiscible consistency model compared to the traditional consistency one. Furthermore, we show the FID is even lower with our method used, confirming the fidelity of our generated images. We also provide images generated from models trained with traditional and immiscible models experiencing the same training steps, where we see those from immiscible models are much more complete and clearer, further proving the training efficiency enhancement resulted from the Immiscible Diffusion. Examples are shown in Fig. 1 right. Deeper analysis shows that our method, though with only one line of code and involving 2̃image-noise datapoint distance changes, achieves all the benefits above in a negligible running time.
@@
  • We identify the issue of the image-noise matching relation, which leads to the slow convergence of diffusion training.
  • We propose a simple and effective method, Immiscible Diffusion, a strategy that only requires one-line of code, to improve training efficiency for diffusion training.
  • Experiments demonstrate the effectiveness of our proposed method on several popular diffusion models across multiple datasets. Additionally, we conduct thorough analyses and ablation studies to elucidate how our method works.
@@ 2.1 Diffusion Model with Efficient Inference
Diffusion models [39, 9, 30, 27] have been attracting huge attention because of its high-fidelity image and video generation [8, 10, 26], its data-efficient perception [40, 24, 44], and even its representation ability for robotics [2, 28, 1]. However, the slow inference is one of the key bottlenecks for diffusion models. To address this issue, various approaches have been proposed. For instance, techniques such as DDIM [35] have reduced the number of denoising steps from 1000 to 10, significantly speeding up the process. Furthermore, the introduction of consistency models [38] and LCM [25], which utilize the properties of a self-consistency, enables denoising in as few as 1-4 steps, further enhancing the generation speed of diffusion models. Subsequently, the development of SD-turbo [33], which leverages GAN [6] loss for high-definition image generation in a single step. The Consistency Trajectory Models [12, 29] improve the generation quality of consistency models, accelerate research on efficient inference for diffusion models. Additionally, beyond reducing denoising steps, efforts to improve inference efficiency of single function evaluation are being explored in various ways, including model quantization [19], and partitioning the generative components [18]. Moreover, StreamDiffusion [13] streamlines denoising steps to achieve real-time inference at the pipeline level optimization. The improvement of the inference efficiency significantly pushes forward the real applications based on Diffusion Models. Yet accelerating diffusion training is still under-explored.
@@ 2.2 Diffusion Model with Efficient Training
Improving the training efficiency of diffusion models is crucial. Various strategies have been proposed, including architectural modifications [11], approximating the diffusion phase with flow [20], and designing parameter choices [37] etc. Specifically, in [11], the authors discover that the magnitude of activation and the magnitude of neural weights significantly impact the training dynamics of diffusion models. They propose adjusting the activation magnitude and standardizing neural weights, as well as modifying the normalization layers to make the diffusion training in a more smooth dynamic. Besides, Song et al. [37] aim to enhance the training efficiency of consistency models through customized design choices, significantly improving both training speed and fidelity. Furthermore, leveraging approximation strategies based on ODE assumptions [20] improves not only inference efficiency but also training efficiency since diffusion trajectories is prone to deterministic. Beyond improving the diffusion training with either adjusting architecture or selecting parameters, Wang et al. [43] introduce a novel patch strategy to control the ease of diffusion training, achieving both training and data efficiency. Concurrently, Wang et al. [42] notice that the denoising of some noisy diffusion steps contain little information and are too easy-to-learn, so focusing more on other steps would significantly enhance the training efficiency. Our method differs from previous works by directly highlighting an under-explored aspect: the relationship between image data and noise, which plays a crucial role in diffusion training. Our proposed Immiscible Diffusion is extremely simple yet significantly improves training efficiency.
@@ 3.1 Physics Intuition
However, we notice that the mixing can also be organized when the sources are immiscible. Under that circumstance, the sources would take different continuous areas after the diffusion, while the whole diffuse-able area remains the same, as shown in the right image of Fig. 2 (a). Thereafter, the reversal process becomes smooth. Inspired by Immiscible Diffusion, we then introduce it to the diffusion models, aiming to make the optimization easier and to achieve a higher training efficiency.
@@ 3.2 Immmiscible Diffusion Model
Similar to the physics phenomenon, we find that for diffusion models, any images are diffused to every corner of the noise space, which also means that each noise point can go back to any image. This would cause the denoise model to be confused on which image to go to, as shown in the left of Fig. 2 (b).
Mimicing the immiscible phenomenon in physics, we hope to design a similar process where each noise point is only matched to limited images, so as to avoid the confusion for the denoise model. However, the noise space must strictly remain Gaussian to help the sampling process. Therefore, we propose Immiscible Diffusion, which assigns the batch of noises to the batch of images during training according to the image-noise distance in their shared space. We minimize the total distance of image-noise pairs in a batch during the assignment. After assignment, the noise is still Gaussian, while each noise is assigned to nearer images like what happens in the immiscible phenomenon, which significantly eases the difficulties for the denoising. Fig. 2 (b) right shows an ideal example of the Immiscible Diffusion, where the noise corresponding to each image is clearly separated.
@@
1:Input: Image batch $x_{b}$, random noise batch $n_{rand,b}$, diffusion steps $t_{b}$ and diffusion schedule $\alpha$
@@
Figure 3: Feature analysis of traditional (miscible) and immiscible DDIM with 20 sampling steps. Note $T=20$ represents the layer denoising the pure noise and $T=1$ represents the last denoising layer before generated image output. We show that while the two noises are similar, the denoised image of immiscible DDIM significantly outperforms that of the traditional one. The reason behind this is traditional methods cannot successfully predict noises at noisy layers. We also notice that the that utilizing only the predicted noise for a single noisy layer, immiscible DDIM can generate an overall reasonable image while traditional DDIM can not.
@@ 3.3 Mathematical Illustration
In this section we mathematically elucidate the denoising difficulty for traditional diffusion models based on DDIM [35] and how our proposed Immiscible Diffusion reduces such difficulties.
In DDIM, we know for any image data-point $x_{0}$, when the diffusion step $t\to\infty$,
$p(x_{t}\mid x_{0})=\mathcal{N}(0,I)=p(x_{t}),\quad where\quad x_{t}=\sqrt{\alpha_{t}}x_{0}+\sqrt{1-\alpha_{t}}\epsilon,\quad\epsilon\sim\mathcal{N}(0,1).$ (1)
Utilizing Bayes’ Rules, we can also interpret Equation 1 as:
$\forall x_{0},x_{t},\quad p(x_{0}\mid x_{t})=\frac{p(x_{t}\mid x_{0})\cdot p(x_{0})}{p(x_{t})}=p(x_{0})\quad when\quad t\to\infty,$ (2)
which indicates that when the model is ideally trained, the distributions of the denoised images for any noise data-point are the same, which is equal to the distribution of the image data.
However, when looking into the last diffusion layer $t$ for our practical training goal $\epsilon_{\theta}(x_{t},t)$, which is the estimated noise at step $t$, we see that for a specific noise point $x_{t}$
$\epsilon_{\theta}(x_{t},t)$ $=\sum_{x_{0}}(ax_{0}-bx_{t})p(x_{0}\mid x_{t})=a\sum_{x_{0}}x_{0}p(x_{0}\mid x_{t})-bx_{t}\sum_{x_{0}}p(x_{0}\mid x_{t})$ (3)
$=a\sum_{x_{0}}x_{0}p(x_{0})-bx_{t}=a\bar{x}_{0}-bx_{t}$
where $a=\frac{1}{\sqrt{1-\alpha_{t}}}$ and $b=\frac{\sqrt{\alpha_{t}}}{\sqrt{1-\alpha_{t}}}$ are constants, and $\bar{x_{0}}$is an average of images in the dataset. When the number of images is large enough, $\bar{x_{0}}\simeq c$ where $c$ is a constant for all pixels. This is especially true when the image dataset is diverse enough, i.e. containing rich image classes. In this way, we find that
$\epsilon_{\theta}(x_{t},t)=ac-bx_{t},$ (4)
which does not provide any meaningful information as it is pointing to a solid color image.
As shown in Fig. 3 (a), we study the predicted noises of different layers in a DDIM model with total inference steps of 20 to illustrate our discovery. Here $T=20$ means the layer of pure noise and $T=1$ is the final denoising layer outputing the generated image. We can see that the predicted noise for DDIM at $T=20$ does not provide much useful information, while subtracting this noise for the image does not provide any distinguishable image, which all support our hypothesis for the difficulty in denoising when $t$ is large. Similar observations are also shown in the concurrent work [42]. However, in our Immiscible Diffusion, while for each batch, we still have
$p(\{x_{t}\}_{batch}\mid\{x_{0}\}_{batch})=\mathcal{N}(0,I),\quad and\quad p(x_{t})=\mathcal{N}(0,I),$ (5)
for each specific $x_{t}$ or $x_{0}$, the conditional noise distribution does not follow the Gaussian distribution because of the batch-wise noise assignment
$p(x_{t}\mid x_{0})\neq\mathcal{N}(0,I).$ (6)
@@
$p(x_{t}\mid x_{0})=f(x_{t}-x_{0})\mathcal{N}(0,I),$ (7)
where $f(x)$ is a function denoting the influence of assignment on the conditional distribution of $x_{t}$. Apparently, according to the definition of linear assignment problem [15], $f(x)$ decreases when $x$ increases its norm, specifically L2 norm as our default setting.
Therefore, from Equation 2 and 7, we have
$\forall x_{0},x_{t},\quad p(x_{0}\mid x_{t})=f(x_{t}-x_{0})p(x_{0}),$ (8)
@@
For noise prediction task, we see that
$\epsilon_{\theta}(x_{t},t)=\sum_{x_{0}}(ax_{0}-bx_{t})p(x_{0}\mid x_{t})=a\sum_{x_{0}}f(x_{t}-x_{0})x_{0}p(x_{0})-bx_{t}=a\overline{x_{0}f(x_{t}-x_{0})}-bx_{t}$ (9)
where $\overline{x_{0}f(x_{t}-x_{0})}$ is the weighted average of $x_{0}$ with more weights on image data-points closer to the noisy data-point $x_{t}$ itself. Therefore, the noise predicted would lead to the average of nearby image data-points, which makes more senses than pointing to a constant. Indeed, in Fig. 3, we see that even for the pure noise layer, immiscible DDIM can predict noise effectively pointing to the shape of the horse image, and the 1-step prediction by substracting the predicted noise shows the outline of the horse correctly.
Note: The assignment problem has been extensively studied for decades [4, 16, 5]. In this paper, we use Hungarian algorithm [16] as our main assignment method. However, Hungarian matching has high complexity with $O(N^{3})$, which drastically slow down the training especially when we have high-dimensional image data (e.g., even using the mini image data $32\times 32\times 3=3072$). To mitigate this issue, we make a novel use of quantization for the image data and noise, i.e., we quantize the $fp32$ image and noise data to $fp16$ or $fp8$ to conduct the assignment, while keeping the same-precision input to diffusion models. This trick significantly reduces the overhead to a negligible level.

### v2

@@ Abstract
In this paper, we point out that suboptimal noise-data mapping leads to slow training of diffusion models. During diffusion training, current methods diffuse each image across the entire noise space, resulting in a mixture of all images at every point in the noise layer. We emphasize that this random mixture of noise-data mapping complicates the optimization of the denoising function in diffusion models. Drawing inspiration from the immiscibility phenomenon in physics, we propose Immiscible Diffusion, a simple and effective method to improve the random mixture of noise-data mapping. In physics, miscibility can vary according to various intermolecular forces. Thus, immiscibility means that the mixing of molecular sources is distinguishable. Inspired by this concept, we propose an assignment-then-diffusion training strategy to achieve Immiscible Diffusion. As one example, prior to diffusing the image data into noise, we assign diffusion target noise for the image data by minimizing the total image-noise pair distance in a mini-batch. The assignment functions analogously to external forces to expel the diffuse-able areas of images, thus mitigating the inherent difficulties in diffusion training. Our approach is remarkably simple, requiring only one line of code to restrict the diffuse-able area for each image while preserving the Gaussian distribution of noise. In this way, each image is preferably projected to nearby noise. To address the high complexity of the assignment algorithm, we employ a quantized assignment strategy, which significantly reduces the computational overhead to a negligible level (e.g. 22.8ms for a large batch size of 1024 on an A6000). Experiments demonstrate that our method can achieve up to 3x faster training for unconditional Consistency Models on the CIFAR dataset, as well as for DDIM and Stable Diffusion on CelebA and ImageNet dataset, and in class-conditional training and fine-tuning. In addition, we conducted a thorough analysis that sheds light on how it improves diffusion training speed while improving fidelity. The code is available at https://yhli123.github.io/immiscible-diffusion
Figure 1: Immiscible Diffusion can use a single line of code to efficiently achieve immiscibility by re-assigning a batch of noise to images. This process results in only a 2% reduction in distance post-assignment, leading to up to 3x increased training efficiency on top of the Consistency Model for CIFAR Dataset. Additionally, Immiscible Diffusion significantly enhances the image quality of Stable Diffusion for both unconditional and conditional generation tasks, and for both training from scratch and fine-tuning training tasks, on ImageNet Dataset within the same number of training steps.
@@ 1 Introduction
Diffusion models have made impressive progress in image generation by framing the process as a phase of denoising random Gaussian noise into the final image. Despite the advancements, training a diffusion model is resource intensive. For example, even in the primary image dataset CIFAR-10, the representative few-step diffusion model, Consistency Model [47], requires training for 10 days on 4 A6000 GPUs to reach a desired FID score of around 10. Similarly, with fewer model parameters, multiple-step diffusion model DDIM [44] still requires 24 hours on an A5000 GPU on the CIFAR-10 dataset. Although recent remarkable achievements in accelerating the inference of diffusion models [19, 33, 47, 28, 30, 31] have been accomplished, the inefficiency of diffusion training remains a significant bottleneck, hindering the iterative development of vision generative AI.
@@
Previous methods for improving diffusion training have focused on various strategies, such as balancing the impact of activation layers and neural weights [16], modifying hyperparameters and design choices [46], and leveraging patchifying strategies [53] etc. Specifically, Karras et al. [16] modifies the activation magnitude, neural weight standardization, and group normalization, achieving significant acceleration in diffusion training. Besides, previous work [46] proposes a customized method for the Consistency Model to improve the performance and diffusion training. Our method is orthogonal to these previous methods. We got inspired by the Immiscible Diffusion in physics. As illustrated in Fig. 2 (a) left, miscible particles tightly jumble together after the diffusion process, making it difficult to separate them individually during the denoising phase. However, when the particles are rendered immiscible, they can still achieve a similar overall distribution while remaining clearly distinguishable (see Fig. 2 (a) right). This insight inspires our strategy for improving the disentanglement of diffused data.
@@
We draw an analogy from the phenomenon of Immiscible Diffusion and relate the distribution of image data to the behavior of particles discussed above. In traditional diffusion processes, each image can be diffused to any point in the noise space, and conversely, each point in the noise space can be denoised to any source image, as illustrated in the left image of Fig. 2 (b). We hypothesize that the jumbled image-noise mapping creates a miscible diffusion effect and makes the optimization of the diffusion model difficult. Inspired by the Immiscible Diffusion, we are motivated to make the mixed diffusion phase distinguishable.
We propose one simple Immiscible Diffusion method. Note that we still sample Gaussian noise but perform a batch-wise assignment of noise to each image based on the distance between them during training. This approach ensures that each image is only diffused to surrounding areas while maintaining the overall Gaussian distribution of all noises. This technique was also used in flow matching-based methods [36, 50] for optimizing the image-noise flow. Nevertheless, we find that the image-noise distance is reduced by only ~2% after assignment, as provided in Part 4.3. This motivates us to ask which factors dominate the performance improvement? To investigate it, we propose another method that qualifies Immiscible Diffusion in Part 4.3. Experiments show that the training efficiency improvement is comparable to the function of flow optimization, demonstrating that immiscibility is the dominant factor.
@@
Figure 2: Physics illustration of Immiscible Diffusion. (a) depict the miscible and Immiscible Diffusion phenomenon in physics, while (b) demonstrate the image-noise pair relation in vanilla (miscible) and Immiscible Diffusion method.
@@
However, technically, to achieve immiscibility with the image data-noise assignment has an O($N^{2}logN$)-O($N^{3}$) complexity. This introduces significant overhead during training, especially for large-scale training with huge batch sizes and high-resolution images. To address this, we employ a novel quantization method during assignment. We quantize the noise and image data into low-precision formats (e.g., 16-bit) during conducting the assignment algorithm. We highlight that this assignment operation only involves one line of code, and is performed only during the training phase without modifying the model architecture, the noise scheduler, the sampler, or the method of inference.
@@
We conduct extensive experiments on three common modes: unconditional, conditional, and fine-tuning on three diffusion baselines: Consistency Models, DDIM and Stable Diffusion and three datasets: CIFAR-10, CelebA and ImageNet datasets. Results show that our proposed method significantly improves the training efficiency in all experiments. Specifically, we achieve 3x training efficiency for the CIFAR-10 dataset with immiscible unconditional Consistency Model compared to the original Consistency Model. Furthermore, we show that the FID is even lower with our method used, confirming the fidelity of our generated images. We also provide images generated from models trained with vanilla and immiscible models experiencing the same training steps, where we see that those from immiscible models are much more complete and clearer, further proving the training efficiency enhancement resulted from the Immiscible Diffusion. Examples are shown on the right of Fig. 1. Deeper analysis shows that our method, although with only one line of code and involving ~2% image-noise datapoint distance changes, achieves all the benefits above in negligible running time.
@@
  • We clearly and specifically identify the miscibility issue in noisy diffusion steps, which leads to slow convergence of diffusion training.
  • To tackle the miscibility issue, we propose a simple and effective method, Immiscible Diffusion, a strategy that can only requires one-line of code, to improve training efficiency for diffusion training.
  • Experiments demonstrate the effectiveness of our proposed method on several popular diffusion models across multiple datasets, and across unconditional, conditional, and fine-tuning tasks. In addition, we conducted thorough analyses and ablation studies to elucidate how our method works and dominates the effectiveness.
@@ 2.1 Diffusion Model with Efficient Inference
Diffusion models [48, 12, 39, 35] have been attracting huge attentions because of their high-fidelity image and video generation [11, 13, 34], data-efficient perception [49, 32, 55], and even representation abilities for robotics [3, 37, 1]. However, slow inference is one of the key bottlenecks for diffusion models. To address this issue, various approaches have been proposed. For instance, techniques such as DDIM [44] have reduced the number of denoising steps from 1000 to 10, significantly speeding up the process. Furthermore, the introduction of Consistency Models [47] and LCM [33], which utilize the properties of self-consistency, enables denoising in as few as 1-4 steps, further enhancing the generation speed of diffusion models. Subsequently, the development of SD-turbo [42], which leverages GAN [8] loss for high-definition image generation in a single step, has occurred. The Consistency Trajectory Models [18, 38] improve the generation quality of Consistency Models and accelerate research on efficient inference for diffusion models. Additionally, beyond reducing denoising steps, efforts to improve the inference efficiency of single function evaluation are being explored in various ways, including model quantization [26] and partitioning the generative components [25]. Moreover, StreamDiffusion [19] streamlines denoising steps to achieve real-time inference at the pipeline level optimization. The improvement of the inference efficiency significantly pushes forward real applications based on diffusion models. Yet accelerating diffusion training is still under-explored.
@@ 2.2 Diffusion Model with Efficient Training
Improving the training efficiency of diffusion models is crucial. Various strategies have been proposed, including architectural modifications [16], approximating the diffusion phase with flow [28], and designing parameter choices [46] etc. Specifically, in [16], the authors discover that the magnitude of activation and the magnitude of neural weights significantly impact the training dynamics of diffusion models. They propose adjusting the activation magnitude and standardizing neural weights, as well as modifying the normalization layers to make diffusion training in a more smooth dynamic. Besides, Song et al. [46] aims to enhance the training efficiency of Consistency Models through customized design choices, significantly improving both training speed and fidelity. Furthermore, leveraging approximation strategies based on ODE assumptions [28] improves not only inference efficiency but also training efficiency since diffusion trajectories are prone to deterministic. Beyond improving diffusion training with either architecture adjustment or selection parameters, Wang et al. [53] introduce a novel patch strategy to control the ease of diffusion training, achieving both training and data efficiency. Gleichzeitig, Wang et al. [52] notices that the denoising of some noisy diffusion steps contains little information and is too easy to learn, so focusing more on other steps would significantly improve training efficiency. Our method differs from previous works by clearly highlighting an under-explored problem: the miscibility problem of image data in the noise space, which plays a crucial role in training diffusion models. Our proposed Immiscible Diffusion is extremely simple yet significantly improves training efficiency.
2.3 Image-noise Optimal Transport in Generative Models
In ODE-based methods such as flow matching [27], straightening the flow with optimal transport (OT) has been used as a tool to improve the generation performance. Specifically, optimizing image-noise transport in a batch [36, 50] has been found to be an effective way to improve performance. Training efficiency improvement was also observed [36, 50], and explained or posited with reduction of the variance of the training goal. However, the standard deviation reduction is only ~4% in [36]. Moving forward, [24] pointed out the curvature problem in the ODE paths caused by the collapse of the reverse trajectories in the average direction. However, they replaced Gaussian noise sampling with a VAE encoder-style structure to eliminate such curvatures, which destroys the strict Gaussian distribution in the noise space. Concurrent to our work, [17] applies batch-wise OT to diffusion models to achieve better FIDs, making posits in curvature reduction for the enhancement. Several methods were proposed to improve the speeds and effectiveness of OT, such as pre-training with PF-ODE [54], using Schrodinger bridging [5], utilizing conditional Wasserstein distance [2] and generator-induced coupling in Consistency Models [14]. However, we are the first to emphasize that the dominant reason for training efficiency enhancement is the miscibility problem in noisy layers, which we prove in Part 4.3. Furthermore, most of previous methods are for unconditional flow matching methods only. Our work demonstrates the effectiveness in multiple diffusion models, datasets, conditional generation, and fine-tuning experiments.
@@ 3.1 Physics Intuition
However, we notice that mixing can also be organized when sources are immiscible. Under that circumstance, the sources would take different continuous areas after diffusion, while the whole diffused area remains the same, as shown in the right image of Fig. 2 (a). Thereafter, the reversal process becomes smooth. Inspired by Immiscible Diffusion, we then introduce it to the diffusion models, with the aim of making the optimization easier and to achieve a higher training efficiency.
@@ 3.2 Immmiscible Diffusion Model
Similar to the physics phenomenon, we find that for diffusion models, any images are diffused to every corner of the noise space, which also means that each noise point can go back to any image. This would cause the denoising model to be confused on which image to go to, as shown in the left of Fig. 2 (b).
Mimicing the immiscible phenomenon in physics, we hope to design similar processes where each noise point is only matched to limited images, so as to avoid the confusion for the denoising model. However, the noise space must remain Gaussian to help the sampling process. Therefore, we propose our first implementation way of Immiscible Diffusion, which assigns the batch of noises to the batch of images during training according to the image-noise distance in their shared space. We minimize the total distance of the image-noise pairs in a batch during assignment. Here we use the L2 distance for assignment, which is ablated in Part A.1.3 in the Supplemental Materials. After assignment, the noise is still Gaussian, while each noise is assigned to nearer images like what happens in the immiscibility phenomenon, which significantly eases the difficulties for the denoising. Fig. 2 (b) right illustrates an extreme example of the Immiscible Diffusion, where the noise corresponding to each image is relatively separated.
@@
1:Input: Image batch $x_{b}$, random noise batch $n_{rand,b}$, sampled diffusion steps $t_{b}$ and diffusion schedule $\alpha$
@@
While linear assignment qualifies Immiscible Diffusion, Immiscible Diffusion can be achieved with multiple paths. In part 4.3, we ablate the linear assignment by letting it remain immiscible while disqualifying optimal transport, proving that immiscibility plays the dominant role in performance improvement.
@@ 3.3 Mathematical Illustration
In this section, we mathematically elucidate the denoising difficulty for traditional diffusion models based on DDIM [44, 12] and how our proposed Immiscible Diffusion reduces such difficulties.
In DDIM, we know for any image data-point $x_{0}$, when it is diffused to the last diffusion step T, i.e. $t=T$, the image is sufficiently wiped out and nearly only gaussian noise is remaining, therefore,
$q(x_{T}\mid x_{0})\approx\mathcal{N}(x_{T};0,I)\approx p(x_{T}),where\quad x_{T}(x_{0},\epsilon)=\sqrt{\overline{\alpha_{t}}}x_{0}+\sqrt{1-\overline{\alpha_{t}}}\epsilon,\quad\epsilon\sim\mathcal{N}(0,1),$ (1)
Where $q$ refers to the Utilizing Bayes’ Rules and Equation 1, we can find that for a specific $x_{T}$:
$\quad p(x_{0}\mid x_{T})=\frac{q(x_{T}\mid x_{0})\cdot p(x_{0})}{p(x_{T})}\approx p(x_{0})\quad,$ (2)
which indicates that the distributions of the corresponding images for any noise data-point are the same as the distribution of all images.
Figure 3: Feature analysis of vanilla (miscible) and immiscible DDIM. Referring to [45], $\tau=S$ represents the layer denoising from the pure noise. We show that while the two sampled noises are similar, the denoised image of immiscible DDIM significantly outperforms that of the traditional one, generating an overall reasonable image. The reason behind this is traditional methods cannot successfully predict noises at noisy layers.
The simplified training objective in DDIM [44, 12] is the added noise $\epsilon(x_{t},t)$ at each diffusion step $t$ and at the point $x_{t}$. However, we find that for a specific point $x_{T}$ in the noise space at diffusion step $T$,
$\epsilon(x_{T},T)$ $=ax_{0}+bx_{T}=\sum_{x_{0}}(ax_{0}+bx_{T})p(x_{0}\mid x_{T})=a\sum_{x_{0}}x_{0}p(x_{0}\mid x_{T})+bx_{T}\sum_{x_{0}}p(x_{0}\mid x_{T})$ (3)
$=a\sum_{x_{0}}x_{0}p(x_{0})+bx_{T}=a\bar{x_{0}}+bx_{t}$
where $a=-\frac{\sqrt{\overline{\alpha_{t}}}}{\sqrt{1-\overline{\alpha_{t}}}}$ and $b=\frac{1}{\sqrt{1-\overline{\alpha_{t}}}}$ are constants, and $\bar{x_{0}}$is an average of images in the dataset. When the number of images is large enough, $\bar{x_{0}}$ contains little meaningful information.
As shown in Fig. 3 (a), we study the predicted noises of different denoising layers in a DDIM model with total inference steps of 20, to illustrate our discovery. Here we refer to [45], specifying the sampling step $\tau=S$ as the layer that denoises the pure noise, which is equal to the diffusing step $t=T$. We can see that the predicted noise for DDIM at $\tau=S$ (equal to $t=T$ in diffusing) does not provide much useful information, while denoising with this "noise" does not provide any distinguishable image, which all support our hypothesis for the difficulty in denoising when $\tau$ in sampling (or $t$ in diffusing) is large. Similar observations are also shown in the concurrent work [52]. However, in our Immiscible Diffusion, while for each batch, we still have
$p(x_{T})=\mathcal{N}(x_{T};0,I),$ (4)
for each specific data point $x_{T}$ or $x_{0}$, the conditional noise distribution does not follow the Gaussian distribution because of the batch-wise noise assignment
$p(x_{T}\mid x_{0})\neq\mathcal{N}(x_{T};0,I).$ (5)
@@
$p(x_{T}\mid x_{0})=f(x_{T}-x_{0},bs,\dots)\mathcal{N}(x_{T};0,I),$ (6)
where $f$ is a function denoting the influence of assignment on the conditional distribution of $x_{T}$, and $bs$ is the training batch size. Apparently, according to the definition of linear assignment problem [21], $f$ decreases when $x_{T}-x_{0}$ increases its norm, specifically the L2 norm as in our default setting.
Therefore, from Equation 2 and 6, we have
$\quad p(x_{0}\mid x_{T})=f(x_{T}-x_{0},...)p(x_{0}),$ (7)
@@
For the noise prediction task, we see that
$\begin{split}\epsilon(x_{T},T)&=\sum_{x_{0}}(ax_{0}+bx_{T})p(x_{0}\mid x_{T})\\ &=a\sum_{x_{0}}f(x_{T}-x_{0},\dots)x_{0}p(x_{0})+bx_{T}\\ &=a\overline{x_{0}f(x_{T}-x_{0},\dots)}+bx_{T}\end{split}$ (8)
where $\overline{x_{0}f(x_{T}-x_{0},...)}$ is the weighted average of $x_{0}$ with more weights on image data-points closer to the noisy data-point $x_{T}$ itself. Therefore, the noise predicted would lead to the average of nearby image data-points, which makes more senses than pointing to a constant. Indeed, in Fig. 3, we see that even for the pure noise layer, immiscible DDIM can predict the noise effectively pointing to the shape of the horse image, and the prediction in one step by subtracting the predicted noise shows the outline of the horse correctly.
3.4 Accelerating Assignment in Immiscible Diffusion
The assignment problem has been studied extensively for decades [6, 22, 7]. In this paper, we use the Hungarian algorithm [22] as our main assignment method. However, Hungarian matching has high complexity with $O(N^{3})$, which drastically slow down the training especially when we have high-dimensional image data (e.g., even using the mini image data $32\times 32\times 3=3072$). To mitigate this issue, we make a novel use of quantization for image data and noise, that is, we quantize the $fp32$ image and noise data to $fp16$ to carry out the assignment, while maintaining the same precision input to diffusion models. This trick significantly reduces the overhead to a negligible level.
To efficiently perform Immiscible Diffusion when running on multiple GPUs, we assign the image-noise distance matrix computation to each process, and then gather them to execute the assignment. This is particularly important as high resolutions and large batch sizes are frequently required in applications.